OpenStudy (anonymous):

Find derivative of -sqrt(y)/sqrt(x) and double derivative of sqrt(x)+sqrt(y)=1

OpenStudy (anonymous):

derivative of \[-\sqrt{y}/{\sqrt{x}}\] and double derivative of \[\sqrt{x}+\sqrt{y}=1\] Than answers to the problems are supposed to be same but it's not happening for me :-(

OpenStudy (anonymous):

derivative of -sqrt(y)/sqrt(x) is [sqrt(x)*(-1/2sqrt(y))*y'-[(-sqrt(y)]*(1/2sqrt(x))] / [sqrt(x)]^2